OpenStudy (anonymous):

3 + a = y is the same as y = 3 + a. This is an example of which algebraic property?

OpenStudy (anonymous):

Not a choice.

sam (.sam.):

thats "commutative property of addition" it means the order changes. for example: 3+4=4+3 x+4+3=3+4+x Just think of it as backwards. 3+4=4+3. 2+3+56=56+3+2 Also, never forget to put commutative prop. "of addition." Because there's commutative prop. of multiplication which is totally different.

OpenStudy (anonymous):

Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

Communitative Property

OpenStudy (phi):

Don't you think it is the symmetric property? See http://www.mathwords.com/s/symmetric_property.htm The question does not change the order of addition.